# SMART: Self-Generating and Self-Validating Multi-Dimensional Assessment for LLMs’ Mathematical Problem Solving

### v1

@@ 4.1 Experimental Setup
4.2 Performance on the SMART-Bench
The performance of the 21 evaluated LLMs on SMART-Bench is detailed in Tab. 1, which presents scores for final answer accuracy alongside metrics for each of the four dimensions.
Dimension specific-task results. Our results indicate that the evaluated models generally demonstrate a strong capacity for problem understanding, with most LLM@Un scores exceeding 90%. This suggests a general proficiency in grasping relevant information and interpreting problem statements. However, significant performance disparities emerge in the reasoning dimension, where ACC@Re scores range widely from 9.05% to 91.45%. A similar divergence is observed in the reflection task with the highest score (78.62%) being nearly nine times greater than the lowest (8.80%). These findings suggest that symbolic reasoning and error reflection capabilities represent critical bottlenecks, particularly for lower-performing models. In contrast, the arithmetic and refinement tasks appear relatively more tractable for most models, with top-performing LLMs achieving near-perfect scores. For instance, o3 attains an ACC@Ar of 98.45%, and DeepSeek-V3 achieves an ACC@R-m of 96.55%, underscoring their strengths in computational and corrective abilities.
Granular insights. Furthermore, the SMART-Bench framework reveals nuanced performance differences among LLMs that final answer accuracy can obscure. For example, while o4-mini and Claude3.7-Sonnet exhibit similar final answer accuracies (93.75% and 93.53%, respectively), o4-mini demonstrates markedly higher proficiency in the reasoning dimension (90.72%) compared to Claude3.7-Sonnet (74.18%). A similar pattern is observed when comparing o4-mini with DeepSeek-V3, demonstrating SMART-Bench’s capacity to reveal fine-grained performance gaps beyond final accuracy metrics. Additionally, consider o4-mini and GPT-4.1: they achieve comparable scores in the understanding dimension and the reflection & refinement dimension. However, o4-mini’s final answer accuracy (93.75%) substantially surpasses that of GPT-4.1 (59.31%). Our framework attributes this disparity primarily to GPT-4.1’s lower reasoning (73.29% vs. o4-mini’s 90.72%) and arithmetic (45.71% vs. o4-mini’s 98.25%) scores. Thus, the SMART framework facilitates a deeper analysis and interpretation of the underlying causes for performance differences.
Data contamination. Theoretically, LLMs should achieve higher accuracy on the arithmetic task compared to the final answer accuracy on the more complex seed questions, as the arithmetic tasks present simplified numerical computations in mathematical notation. However, our results reveal instances where models such as Gemma3-27B, GLM4-32B, GPT-4o, and GPT-4.1 exhibit lower ACC@Ar scores than their ACC@Fi scores. This counter-intuitive finding may suggest the presence of data contamination in the training sets of these models, where they might have memorized answers to full problems similar to those in standard benchmarks but struggle with the decontextualized arithmetic components presented in our SMART-Bench.
@@ 4.2 Performance on the SMART-Bench
4.3 How Does Task Difficulty Impact Different Dimensions of SMART?
To examine factors affecting model performance across each SMART dimension, we construct new sets of dimension-specific questions with varying difficulty levels and evaluate five leading closed-source LLMs on this dynamic test set. Details of the difficulty design are provided in Appendix D.3.
Fig. 3 presents performance trends across difficulty levels for each SMART dimension. As expected, increased task complexity generally leads to substantial performance degradation, validating the effectiveness of our difficulty calibration. Notably, GPT-4.1 and Claude3.7-Sonnet exhibit heightened sensitivity in the reasoning dimension, with ACC@Re scores dropping sharply from  90% to below 60% as reasoning steps increase. In contrast, the remaining models sustain ACC@Re scores above 80% even beyond six reasoning steps. In the arithmetic dimension, o4-mini maintains robust performance (ACC@Ar > 80%) even with nine-digit numbers (refers to the number of digits in scientific notation instead of the count of operands), whereas GPT-4.1’s accuracy drops below 10%. For reflection tasks, introducing only two error types into the CoT results in a sharp decline in detection accuracy across all models. None can reliably detect all errors when four or more distinct mistake types are present.
Model Perturbation Final Answer Understanding Reasoning Arithmetic
ACC@An PD LLM@Un PD ACC@Reason PD ACC@Ar PD
GPT-4.1 Seed question 59.31 / 95.41 / 73.29 / 45.71 /
+ Noise insertion 41.23 18.08$\downarrow$ 88.52 6.89$\downarrow$ 46.54 26.75$\downarrow$ 39.40 6.31$\downarrow$
+ Adding operation 28.45 30.86$\downarrow$ 91.32 4.09$\downarrow$ 57.53 15.76$\downarrow$ 20.64 25.07$\downarrow$
+ Numerical variation 24.81 34.50$\downarrow$ 92.83 2.58$\downarrow$ 52.81 20.48$\downarrow$ 18.34 27.37$\downarrow$
Claude-3.7 Seed question 93.53 / 95.09 / 74.18 / 98.11 /
+ Noise insertion 87.60 5.93$\downarrow$ 89.87 5.22$\downarrow$ 65.81 8.37$\downarrow$ 91.82 6.29$\downarrow$
+ Adding operation 76.21 17.32$\downarrow$ 92.34 2.75$\downarrow$ 54.38 19.80$\downarrow$ 67.43 24.39$\downarrow$
+ Numerical variation 62.74 30.79$\downarrow$ 93.17 1.92$\downarrow$ 58.92 15.26$\downarrow$ 58.29 39.82$\downarrow$
@@ 4.3 How Does Task Difficulty Impact Different Dimensions of SMART?
Table 2: The performance degradation of evaluation dimensions when three types of perturbations are added to the seed questions. PD refers to the performance drop.
@@
4.4 How Do Fine-grained Dimensions Influence the Performance of Final Answer Accuracy?
@@
Prior work has demonstrated that LLMs exhibit significant performance degradation in final answer accuracy when evaluated on question variations generated through perturbations Li et al. (2024b); Zhu et al. (2023); Li et al. (2024a). However, the specific contributions of different cognitive dimensions to this degradation remain underexplored. Therefore, we adapt three perturbation strategies from Li et al. (2024b) applying them to both the seed questions and their corresponding dimension-specific variants. This allows us to investigate which underlying capabilities are most influential in the observed drop in final answer accuracy under perturbation.
@@
Tab. 2 shows that all evaluated dimensions exhibit noticeable performance drops (PD) under perturbations. When noise is introduced, the reasoning dimension is most affected for both GPT-4.1 (26.75%) and Claude 3.7-Sonnet (8.37%). In contrast, perturbations involving added operations or numerical modifications result in the largest performance drops in the arithmetic dimension for both models. These findings suggest that irrelevant information primarily disrupts reasoning capabilities, while alterations to operations or numerical values predominantly impair arithmetic proficiency. Such fine-grained dimension performance drop ultimately contribute to reduced final answer accuracy.
Figure 4: The confusion matrix of final answer and other dimensions. P means Positive, and N means Negative.
4.5 Is the Final Answer Accuracy Reliable for Measuring Mathematical Capability?
Given the impressive performance of LLMs and the potential for data contamination, there is a concern that models might solve problems correctly without possessing genuine underlying mathematical capability. To investigate this, we conduct experiments to compute confusion matrices comparing final answer correctness with performance on our SMART dimensions, as illustrated in Fig. 4. We posit that true mathematical capability is more accurately reflected by instances where LLMs correctly solve not only the original seed question but also simultaneously succeed in the corresponding reasoning and arithmetic dimension tasks.
Across all confusion matrices, the False Negative (FN) values are consistently non-zero. This indicates that LLMs can sometimes arrive at correct final answers through heuristic shortcuts or other opaque mechanisms, even when their intermediate reasoning or calculation processes are flawed. For example, GPT-4.1 exhibits a notable FN rate of 11.90% in final answer and arithmetic confusion matrix. Similarly, Claude3.7-Sonnet shows an FN rate of 22.95% in final answer and reasoning confusion matrix. These FN cases represent instances where final answer accuracy overestimates the model’s grasp of the intermediate steps.
Conversely, False Positive (FP) scores denote cases where a model successfully completes an intermediate task but ultimately yields an incorrect final answer. Except for GPT-4.1, most evaluated LLMs exhibit relatively low FP rates across various confusion matrices, indicating that accurate intermediate reasoning generally correlates with correct final outputs.
True Positive (TP) scores capture instances where a model not only produces the correct final answer but also performs intermediate reasoning and arithmetic correctly. We regard this TP metric as a more reliable indicator of genuine mathematical problem-solving ability. For high-performing models such as o4-mini, DeepSeek-R1, and Gemini2.5-Pro-Preview, TP scores in the reasoning & arithmetic confusion matrix closely match their ACC@Fi values in Tab. 1. In contrast, GPT-4.1 and Claude3.7-Sonnet exhibit significantly lower TP scores relative to their ACC@Fi, suggesting that their final answer accuracy may overestimate their true reasoning capabilities.
5 Conclusion
@@
We introduce the SMART evaluation framework and its accompanying benchmark, SMART-Bench, for systematic assessment of LLMs in mathematical problem solving. SMART decomposes this complex ability into four interpretable dimensions, Understanding, Reasoning, Arithmetic, and Reflection & Refinement, enabling more granular evaluation than traditional answer-based metrics. A key feature is its self-generating and self-validating mechanism, which iteratively verifies the correctness of generated test data to ensure reliability and scalability. Experiments on 21 open- and closed-source LLMs reveal that reasoning and reflection remain primary bottlenecks. Moreover, our analysis highlights the limitations of final-answer accuracy as a sole indicator and introduces a new metric to better capture true model capability. We hope SMART serves as a foundation for more comprehensive and interpretable evaluations.
6 Limitation
While our proposed framework provides a comprehensive evaluation platform, it is important to acknowledge its inherent limitations. In particular, although Z3 and SMT-LIB effectively handle linear, integer, and some nonlinear constraints, their problem-solving capabilities are restricted. They are unsuitable for highly complex nonlinear problems and certain NP-complete combinatorial tasks. To overcome these limitations, future work will investigate using Lean Moura and Ullrich (2021) to formalize and prove complex mathematical theorems involving higher-order logic and intricate proof structures beyond SMT solvers’ scope.
Appendix A Polya’s Problem-solving Theory
@@
Polya first propose the four step problem solving method in the book ’How to solve it’ Pólya and Conway (1957) to systematically answer questions. According to his theory, it breaks down the process of problem-solving into four steps (Fig.  1): 1) understanding the problem. 2) make a plan. 3) carry out the plan. 4) look back and review.
@@
Appendix B SMART-Bench Generation
B.1 Seed Question Dataset of SMART
The foundation of the SMART benchmark is a seed dataset comprising 2000 problem instances. These were curated from seven widely-used mathematical problem datasets: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), AQuA Ling et al. (2017), MAWPS Koncel-Kedziorski et al. (2016), MATH Hendrycks et al. (2021), and problems from the AIME 2024 competition Huggingface (2024). These initial 2,000 seed questions, along with their subsequently generated dimension-specific variations (four per seed question), form the complete SMART-Bench, totaling 10,000 test instances.
@@ B.1 Seed Question Dataset of SMART
B.2 Dimension-Specific Testing Data and Task Description
@@
B.2.1 Understanding
To generate ground truth for the understanding dimension, we utilize GPT-4o to perform context extraction. The extracted context is structured into the following components: Problem Scenario (describing the overall context of the problem), Goal (specifying what needs to be solved), Known Quantities (listing explicitly provided numerical values or facts), Unknown Quantities (identifying variables or values to be determined), Relationships and Constraints (detailing connections and limitations between pieces of information), and Irrelevant Information (pinpointing details not pertinent to the solution). The prompt employed to guide GPT-4o in extracting this contextual information for ground truth generation is presented in Fig. 14. To enhance the performance of GPT-4o in generating these contexts, we employed a three-shot prompting strategy. To ensure the quality of this generated ground truth, we adopt the human verification and iterative regeneration strategy outlined by Chen et al. (2024). Specifically, after the initial generation of ground truth contexts, a random 10% sample of these is manually assessed by human annotators. If the annotators find less than 80% of the sampled contexts satisfactory, the entire set of contexts is regenerated using a refined approach, and the verification process is repeated.
B.2.2 Reasoning
@@
B.2.3 Arithmetic
The arithmetic ability of LLMs is measured by their performance on pure numeric calculation problems that share the same reasoning logic and final answer as the seed question. Directly converting a seed question into an arithmetic notation problem is challenging for LLMs. This difficulty arises because it requires simplifying complex natural language into structured mathematical operations while preserving the logical relationships between variables. Such a transformation is non-trivial, as the model must accurately interpret the problem’s intent, handle ambiguous phrasing, and correctly map linguistic constructs to arithmetic operations. Therefore, to address this, we first generate an SMT-LIB representation of the seed question, which simplifies the reasoning logic among variables. Subsequently, we utilize GPT-4o to convert this SMT-LIB representation into the arithmetic notation problem. The prompt for that process is shown in Fig.  7 and Fig.  8.
B.2.4 Reflection & Refinement
For the Reflection & Refinement dimension, we first generate Chain-of-Thought (CoT) solutions containing deliberate errors. To create these erroneous CoTs, one of the six error types defined in Fig. 18 (e.g., arithmetic number error, skipped step) is uniformly sampled and injected into a correct CoT. The ground truth for the reflection sub-task is then the specific name of the introduced error type and its order number. For the subsequent refinement sub-task, direct verification of the LLM-corrected CoT is complex. Instead, we evaluate the refined CoT by using GPT-4o to generate a final answer based upon it. The LLM is considered to have successfully passed the refinement task if this GPT-4o-generated answer matches the ground truth of the original seed question.
@@ B.2.4 Reflection & Refinement
Appendix C SMART-Bench Verification
We adopt the neuro-symbolic system to automatically verify the correctness of generated SMT-LIB and its arithmetic notation formed question. This self-validating mechanism can also be use to verify the augmented version of questions which are generated based on the number of operator modification rewriting rules. The process is as following:
  1. 1.
    Initial SMT-LIB Conversion and Validation: We first attempt to convert the original seed question  into an SMT-LIB formula. We employ an SMT solver to compute the solution from this SMT-LIB formula. If the solver-derived result does not match the known ground-truth answer of the original seed problem, the SMT-LIB generation process is iteratively refined and re-verified until consistency is achieved. This ensures a validated symbolic representation of the original problem.
  2. 2.
    Augmentation in SMT-LIB Space: Once a validated SMT-LIB formula for the original problem  is obtained, numerical or operational changes dictated by our rewriting rules are applied directly within this symbolic SMT-LIB space.
  3. 3.
    Automated Ground-Truth Generation for Augmented Problems: The modified SMT-LIB formula, now representing an augmented version of the problem, is then solved using the Z3-solver. The result provided by the solver serves as the automatically generated and verified ground-truth answer for this new, dynamically created problem instance.
@@ Appendix C SMART-Bench Verification
Appendix D Experiments
D.1 Experiment Setting
@@
D.2 Prompt and Evaluation metrics in SMART
D.2.1 Understanding
@@
We employed an LLM-as-a-judge method Zheng et al. (2023) to evaluate the quality of generated context. In this method, GPT-4o is used impartially to assess semantic similarity between the background context extracted by the evaluated LLM and the reference ground-truth, assigning a score from 1 to 100.
@@
D.2.2 Reasoning
For the reasoning dimension, we introduce a symbolic formalization task to evaluate the symbolic reasoning capability of LLMs. The input for this task is the seed question, and LLMs are asked to generate the SMT-LIB expression of the question, without solving the problem. The prompt for this task is shown in Fig. 7. Subsequently, the Z3 solver is used to compute the results of the generated SMT-LIB expression. The evaluation metric for this task is the accuracy of the SMT-LIB formula results. Specifically, the accuracy of generating correct SMT-LIB expressions, determined by the number of correct SMT-LIB (verified by Z3-solver) divided by the total number of problems.
D.2.3 Arithmetic
For the arithmetic dimension, we introduce a numeric calculation task to assess the arithmetic ability of LLMs. The input for this task is an arithmetic notation problem, and LLMs are asked to solve it using the prompt shown in Fig. 9. The arithmetic problem is designed to have the same answer and reasoning logic as its corresponding seed question. The evaluation metrics for this task is accuracy, defined as the number of arithmetic problems solved correctly divided by the total number of problems.
D.2.4 Reflection & Refinement
@@
The evaluation metrics for the reflection task is accuracy. The accuracy in identifying all errors in the chain-of-thought (CoT) solution, computed as the number of problems where all CoT errors are correctly identified divided by the total number of problems. The accuracy of refining the CoT after error identification, measured as the number of problems correctly answered based on the refined CoT divided by the number of problems with correctly identified CoT errors. The combined accuracy for both reflection and refinement tasks, calculated as the number of problems correctly answered after CoT refinement divided by the total number of problems.
D.3 Difficulty setting for Dimension-specific Task
@@
For the reasoning dimension, question complexity is defined by the number of distinct mathematical operations (e.g., +,,×,÷$+,-,\times,\div,mod$) required to formulate the solution. Problems are then categorized into multiple difficulty levels based on this operational count.
@@
D.4 Examples for question with different difficulty setting
@@
D.5 Zero-shot vs. three-shot

### v3

@@ 3. $\bullet$
Human verification
The generation of arithmetic notation problems from SMT-LIB representations, as well as the extraction of structured information as ground truth for the Understanding dimension, are both performed using GPT-4.1 and cannot be validated by the neuro-symbolic system. Thus, we follow the human verification protocol proposed by Chen et al. (2024) to ensure the reliability of these generated data. Specifically, a randomly selected 10% subset of the data is manually reviewed by human annotators. If this sample does not meet quality standards, the entire set is regenerated and re-verified until it passes inspection.
As a result, SMART-Bench comprises 10,000 test instances, including 2,000 original seed questions and 8,000 carefully curated, task-specific variants. Through a combination of neuro-symbolic and human verification, we ensure that each instance meets quality standards.
Experiments
Models
We evaluate 21 recent open- and closed-source LLMs using our SMART framework, covering both general-purpose and reasoning-specialized models. To ensure deterministic outputs, the temperature is set to 0. For each dimension-specific task, we employ a three-shot prompting strategy.
@@
Evaluation Metrics
We employ a combination of evaluators for SMART, including LLM-as-a-Judge (Zheng et al. 2023), the Z3 solver, and rule matching, as summarized in Tab. 1. Additionally, we introduce a new metric, All-Pass Score, which reflects a model’s overall performance across the full problem-solving process.
For the Arithmetic, Reflection and Refinement tasks, we employ a rule-based approach to extract the final answer from the model output and compare it against the ground truth to compute the accuracy-based metric. For example, the accuracy for the Arithmetic task (denoted as ACC@Ar) is computed as follows:
$ACC@Ar$ $=$ $\frac{N_{correct\_answers\_sample}}{N_{total\_sample}}$ (1)
Metrics such as ACC@An (Final answer accuracy), ACC@Re (Reasoning task), ACC@R-t (Reflection task), ACC@R-m (Refinement task), and ACC@RR (R&R task) are computed using the same formulation. For Reasoning task, we do not evaluate the correctness of the generated SMT-LIB expressions, since multiple logically equivalent solutions can exist for a problem. Instead, we validate these symbolic expressions by executing them with the Z3 solver and comparing the solver’s output with the ground-truth answer from the original seed question. For the Refinement task, we GPT-4.1 to generate a final answer based upon the refined COT. The the refined COT is correct if the GPT-4.1-generated response matches the ground truth.
For the Understanding dimension, we adopt the LLM-as-a-Judge evaluation approach, introducing the metric LLM@Un to assess the quality of generated structured contexts. To mitigate potential preference bias (Li et al. ) inherent in LLM-based judging, we independently employ GPT-4.1 and DeepSeek-V3 as two judging models. Each judge evaluates the semantic similarity between the model-generated context and the reference ground truth, assigning a similarity score ranging from 1 to 100. The final LLM@Un score is computed by averaging the two judges’ scores:
$LLM@Un$ $=$ $\frac{LLM@Un_{GPT}+LLM@Un_{DS}}{{2}}$ (2)
The All-Pass Score is an integrated metric (ACC@ALL) that combines performance across all evaluation dimensions. Specifically, a model achieves an All-Pass success if it simultaneously meets the following criteria: (1) obtaining a score of at least 90% on the Understanding task; (2) correctly solving the Reasoning task; (3) correctly solving the Arithmetic task; and (4) successfully completing the entire R&R task. This stringent criterion ensures that the All-Pass Score accurately reflects comprehensive, end-to-end mathematical problem-solving capability.
Model
Perturbation Final Answer Understanding Reasoning Arithmetic R&R
ACC@An PD LLM@Un PD ACC@Reason PD ACC@Ar PD ACC@$R$ PD
GPT-4.1
Seed question 59.31 / 95.41 / 73.29 / 45.71 / 59.90 /
[…]
Claude-3.7
Seed question 93.53 / 95.09 / 74.18 / 98.11 / 57.43 /
[…]
@@
Table 3: The performance degradation of evaluation dimensions when three types of perturbations are added to the seed questions. PD refers to the performance drop. The most affected dimension in each case is highlighted in bold.
Method
Refined
Dimension
ACC@An
ACC@An
with self-refine
Llama3.1-8B Reasoning 45.15 43.85 ($\downarrow 1.30$)
Mistral-Small-24B Reasoning 77.55 78.15 ($\uparrow 0.61$)
GLM-24B Arithmetic 54.13 56.28 ($\uparrow 2.15$)
Qwen2.5-72B Arithmetic 41.05 52.82 ($\uparrow 11.77$)
Table 4: Self-Refine prompting on specific dimension.
Performance on the SMART-Bench
The performance of the 21 evaluated LLMs on SMART-Bench is detailed in Tab. 2, which reports the scores for all evaluation dimensions and final answer accuracy.
Dimension specific-task results
Our results indicate that the LLMs generally demonstrate a strong capacity for problem understanding, with most LLM@Un scores exceeding 90%. This suggests a general proficiency in grasping relevant information and interpreting problem statements. However, significant performance disparities emerge in the Reasoning dimension, where ACC@Re scores range widely from 9.05% to 93.61%. A similar divergence is observed in the Reflection task with the highest score (78.62%) being nearly nine times greater than the lowest (8.80%). These findings suggest that symbolic reasoning and error reflection capabilities represent critical bottlenecks, particularly for smaller models. In contrast, the Arithmetic and Refinement tasks appear relatively less challenging, with leading LLMs achieving near-perfect performance. For example, o3 attains an ACC@Ar of 98.45%, while DeepSeek-V3 reaches an ACC@R-m of 96.55%, demonstrating their strength in computational and corrective abilities.
Granular insights
The SMART-Bench framework uncovers nuanced performance differences among LLMs that are obscured by final answer metrics alone. For example, while o4-mini and Claude3.7-Sonnet exhibit similar final answer accuracies (93.75% and 93.53%, respectively), o4-mini demonstrates markedly higher proficiency in the Reasoning dimension (90.72%) compared to Claude3.7-Sonnet (74.18%). A similar trend is observed when comparing o4-mini and DeepSeek-V3, further illustrating SMART-Bench’s ability to reveal fine-grained gaps that traditional outcome-based metrics miss. Additionally, although o4-mini and GPT-4.1 perform similarly on the Understanding and Reflection & Refinement dimensions, o4-mini’s final answer accuracy (93.75%) is markedly higher than GPT-4.1’s (59.31%). Our framework attributes this disparity primarily to GPT-4.1’s lower Reasoning (73.29% vs. o4-mini’s 90.72%) and Arithmetic (45.71% vs. o4-mini’s 98.25%) scores. Thus, the SMART framework facilitates a deeper analysis and interpretation of the underlying causes for performance differences.
All-Pass Score remains a challenge
To provide a more challenging and discriminative evaluation, we introduce the All-Pass Score, which measures a model’s ability to simultaneously satisfy the success criteria across all dimensions for each seed question. The All-Pass Score effectively distinguishes model performance. The highest All-Pass Score achieved is only 64.87% for o3, which is substantially lower than its final answer accuracy of 91.44%. This metric offers a much clearer assessment of a model’s mathematical ability and highlights considerable room for further improvement.
How Does Task Difficulty Impact Different Dimensions of SMART?
To investigate how task difficulty affects model performance across different SMART dimensions, we construct new sets of dimension-specific questions with varied difficulty levels and evaluate five leading closed-source LLMs on this dynamic test set. Task difficulty is manipulated in the following ways: for the Understanding dimension, by varying the number of added irrelevant sentences; for the Reasoning dimension, by grouping questions according to the number of required reasoning steps; for the Arithmetic dimension, by changing the number of digits (referring to digit length in scientific notation, not the number of operands); and for the R&R dimension, by altering the number of mistakes introduced into the CoT.
It is important to note that modifying the number of digits in arithmetic questions changes the ground-truth answer. To ensure correctness, we simultaneously update both the numerical values in the arithmetic questions and their corresponding SMT-LIB representations, subsequently employing the Z3 solver to generate new ground-truth answers. For the other dimensions, the ground-truth answers remain unchanged. Comprehensive details regarding the difficulty settings are provided in the Appendix.
As shown in Fig. 3, increasing task complexity generally leads to notable performance degradation across all dimensions. Notably, GPT-4.1 and Claude3.7-Sonnet show pronounced sensitivity in the Reasoning dimension, with ACC@Re scores dropping sharply from approximately 90% to below 60% as the number of reasoning steps increases. In contrast, the remaining models maintain ACC@Re scores above 80% even with more than six reasoning steps. In the Arithmetic dimension, o4-mini demonstrates robust performance (ACC@Ar >> 80%) even with nine-digit numbers, whereas GPT-4.1’s accuracy falls below 10%. For Reflection tasks, introducing just two error types into the CoT results in a steep decline in detection accuracy for all models, with none able to reliably detect all errors when four or more distinct mistake types are present.
@@
How Do Fine-grained Dimensions Influence the Performance of Final Answer Accuracy?
@@
Prior work has shown that LLMs experience significant drops in final answer accuracy when evaluated on perturbed versions of questions (Li et al. 2024b; Zhu et al. 2023; Li et al. 2024a). However, the underlying reasons for this degradation remain insufficiently explored. To address this gap, we adapt three perturbation strategies from (Li et al. 2024b) and apply them to both the seed questions and their corresponding dimension-specific variants, aiming to identify which dimensions are most susceptible to performance loss under these perturbations.
@@
As shown in Table 3, all evaluated dimensions exhibit substantial performance drops (PD) under perturbations. When noise is introduced, the reasoning dimension is most affected for both GPT-4.1 (26.75%) and Claude 3.7-Sonnet (8.37%). Conversely, additional operations or numerical modifications lead to the greatest drops in the Arithmetic dimension. These results suggest that irrelevant information primarily undermines reasoning abilities, while changes to operations or numeric values predominantly impact arithmetic proficiency. Ultimately, vulnerabilities across all dimensions collectively reduce final answer accuracy.
Improving LLMs via Self-Refine Prompting on Weak Dimensions
To enhance the mathematical capabilities of LLMs, we apply self-refinement prompting specifically to the weakest step identified in the problem-solving process. The specific prompts are provided in the Appendix. As shown in Tab.4, targeting the reasoning or arithmetic dimension leads to notable performance gains for Gemma3-27B, Mistral-Small, and Qwen2.5-72B. In contrast, Llama3.1-8B experiences a slight performance drop, likely due to its limited capacity for self-reflection. These results demonstrate that the SMART framework is an effective diagnostic tool for pinpointing a model’s weakest dimension, and that targeted intervention on this dimension can improve mathematical performance.
Conclusion
@@
We present the SMART evaluation framework and its corresponding benchmark, SMART-Bench, for the systematic assessment of LLMs in mathematical problem solving. SMART decomposes this complex ability into four fine-grained dimensions—Understanding, Reasoning, Arithmetic, and Reflection & Refinement—and introduces a novel All-Pass Score metric, enabling more comprehensive evaluation. We propose a self-generating and self-validating mechanism that iteratively verifies the correctness of generated test data to construct SMART-Bench. Experiments on 21 open- and closed-source LLMs reveal that Reasoning and Reflection remain the primary bottlenecks, and that targeted improvements to weak dimensions can enhance the mathematical capabilities of LLMs. We hope that SMART provides a foundation for more comprehensive and interpretable evaluation of large language models in the future.
Polya’s Problem-solving Theory
@@
Polya first propose the four step problem solving method in the book ’How to solve it’ (Pólya and Conway 1957) to systematically answer questions. According to his theory, it breaks down the process of problem-solving into four steps (Fig.  4): 1) understanding the problem. 2) make a plan. 3) carry out the plan. 4) look back and review.
@@
Figure 4: Illustration of the SMART framework for evaluating LLMs. According to Polya’s theory, humans solve mathematical problems through four steps: understanding the problem, devising a plan, carrying out the plan, and looking back. Based on these steps, we define four evaluation dimensions and corresponding tasks to systematically assess LLMs.
SMART-Bench Generation
Seed Question Dataset of SMART
The foundation of the SMART benchmark is a seed dataset comprising 2000 problem instances. These were curated from seven widely-used mathematical problem datasets: GSM8K (Cobbe et al. 2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), ASDiv (Miao, Liang, and Su 2020), AQuA (Ling et al. 2017), MAWPS (Koncel-Kedziorski et al. 2016), MATH (Hendrycks et al. 2021), and problems from the AIME 2024 competition (Huggingface 2024). These initial 2,000 seed questions, along with their subsequently generated dimension-specific variations (four per seed question), form the complete SMART-Bench, totaling 10,000 test instances.
@@
Dimension-Specific Testing Data and Task Description
@@
Understanding
To generate ground truth for the understanding dimension, we utilize GPT-4.1 to perform context extraction. The extracted context is structured into the following components: Problem Scenario (describing the overall context of the problem), Goal (specifying what needs to be solved), Known Quantities (listing explicitly provided numerical values or facts), Unknown Quantities (identifying variables or values to be determined), Relationships and Constraints (detailing connections and limitations between pieces of information), and Irrelevant Information (pinpointing details not pertinent to the solution). The prompt employed to guide GPT-4.1 in extracting this contextual information for ground truth generation is presented in Fig. 14.
Reasoning
@@
Arithmetic
The arithmetic ability of LLMs is measured by their performance on pure numeric calculation problems that share the same reasoning logic and final answer as the seed question. Directly converting a seed question into an arithmetic notation problem is challenging for LLMs. This difficulty arises because it requires simplifying complex natural language into structured mathematical operations while preserving the logical relationships between variables. Such a transformation is non-trivial, as the model must accurately interpret the problem’s intent, handle ambiguous phrasing, and correctly map linguistic constructs to arithmetic operations. Therefore, to address this, we first generate an SMT-LIB representation of the seed question, which simplifies the reasoning logic among variables. Subsequently, we utilize GPT-4.1 to convert this SMT-LIB representation into the arithmetic notation problem and then manually checked by human annotators. The prompt for that process is shown in Fig.  7 and Fig.  8.
Reflection & Refinement
For the Reflection & Refinement dimension, we first generate Chain-of-Thought (CoT) solutions containing deliberate errors. To create these erroneous CoTs, one of the six error types defined in Fig. 20 (e.g., arithmetic number error, skipped step) is uniformly sampled and injected into a correct CoT. The ground truth for the reflection sub-task is then the specific name of the introduced error type and its order number. For the subsequent refinement sub-task, direct verification of the LLM-corrected CoT is complex. Instead, we evaluate the refined CoT by using GPT-4.1 to generate a final answer based upon it. The LLM is considered to have successfully passed the refinement task if this GPT-4.1-generated answer matches the ground truth of the original seed question.
@@
Limitation
While our proposed framework provides a comprehensive evaluation platform, it is important to acknowledge its inherent limitations. In particular, although Z3 and SMT-LIB effectively handle linear, integer, and some nonlinear constraints, their problem-solving capabilities are restricted. They are unsuitable for highly complex nonlinear problems and certain NP-complete combinatorial tasks. To overcome these limitations, future work will investigate using Lean (Moura and Ullrich 2021) to formalize and prove complex mathematical theorems involving higher-order logic and intricate proof structures beyond SMT solvers’ scope.
Experiments
Experiment Setting
We evaluate 21 recent open-source and closed-source LLMs using our SMART evaluation framework. The open-source models include Phi4 (Abdin et al. 2024), Gemma3 (Team et al. 2025), GLM4 (GLM et al. 2024), Qwen2.5 (Team 2024), Qwen3 (Yang et al. 2025), Llama3 (AI@Meta 2024), Llama4 (Meta 2025), and Mistral (MistralAITeam 2024). The closed-source models assessed are GPT-4o (Achiam et al. 2023), GPT-4.1 (OpenAI 2025a), o4-mini, o3 (OpenAI 2025b), DeepSeek-V3 (Liu et al. 2024), DeepSeek-R1 (Guo et al. 2025), Claude3.5 (anthropic 2024), Claude3.7 (anthropic 2025), and Gemini2.5 (Google 2024).
@@
Prompts for Evaluation in SMART
Understanding
@@
Reasoning
For the reasoning dimension, we introduce a symbolic formalization task to evaluate the symbolic reasoning capability of LLMs. The input for this task is the seed question, and LLMs are asked to generate the SMT-LIB expression of the question, without solving the problem. The prompt for this task is shown in Fig. 7. Subsequently, the Z3 solver is used to compute the results of the generated SMT-LIB expression.
Arithmetic
For the arithmetic dimension, we introduce a numeric calculation task to assess the arithmetic ability of LLMs. The input for this task is an arithmetic notation problem, and LLMs are asked to solve it using the prompt shown in Fig. 9. The arithmetic problem is designed to have the same answer and reasoning logic as its corresponding seed question.
Reflection & Refinement
@@
The Self-Refine Prompting
We apply self-refinement prompting specifically to the weakest step identified in the problem-solving process to improve the mathematic capability of LLMs. The prompts used for self-refinement on the reasoning and arithmetic dimensions are shown in Fig.15 and Fig.16.
Figure 15: The prompt for LLMs to self-refine the Reasoning dimension.
Figure 16: The prompt for LLMs to self-refine the Arithmetic dimension.
Difficulty setting for Dimension-specific Task
@@
For the reasoning dimension, question complexity is defined by the number of distinct mathematical operations (e.g., $\times,\div,mod$) required to formulate the solution. Problems are then categorized into multiple difficulty levels based on this operational count.
@@
Examples for question with different difficulty setting
@@
Zero-shot vs. three-shot